OpenStudy (anonymous):

3^4 + 4 ⋅ 5 = ____. (Input whole numbers only.) Numerical Answers Expected! Answer for Blank 1:

OpenStudy (anonymous):

What's 3^4?

OpenStudy (anonymous):

\(3^4 \rightarrow 3 \times 3 \times 3 \times 3\)

OpenStudy (ribhu):

101

OpenStudy (anonymous):

Besides, 101 is incorrect.

OpenStudy (anonymous):

lol i was not giving that dude the answer i was just choosing a raindom one lol

OpenStudy (anonymous):

9x3x3

OpenStudy (anonymous):

What's 3 * 3 * 3 * 3?

OpenStudy (anonymous):

27x3

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Now do it one more time.

OpenStudy (ribhu):

101 is not wrong

OpenStudy (anonymous):

81

OpenStudy (anonymous):

Yes, now add 81 + 4.

OpenStudy (anonymous):

no its 4 x 5 the rules of pemds remember

OpenStudy (anonymous):

Oh right..lol.

OpenStudy (anonymous):

Good catch, what's 4 * 5?

OpenStudy (ribhu):

it is 81 + 20 =101

OpenStudy (anonymous):

20 + 81 = 101

OpenStudy (anonymous):

Yes, you got it.

OpenStudy (ribhu):

now whats the answer u @iGreen.

OpenStudy (anonymous):

@ganeshie8 @shrutipande9

OpenStudy (anonymous):

@ribhu Stop giving direct answers, I've already told you enough times.

OpenStudy (anonymous):

lol i beat the master at his game i caught you. lol none of us are perfect

OpenStudy (anonymous):

I know :)

OpenStudy (ribhu):

buddy it just needed a number by the way